What volume of of 0.25 M H2SO4 is needed to react with 40.0 ml of 0.25 M Barium hydroxide?
1
Expert's answer
2020-10-02T14:22:25-0400

H2SO4+Ba(OH)2=BaSO4+H2O

"C1\times V1=C2\times V2"

"V2=\frac{C2\times V2}{C1}=\frac{0.04\times0.25}{0.25}=0.04 L"